Help marginal revenue for the production of the gven number of bushels (a) The (Round訇four decimal places as needed) The marginal revenue is hundired million doilars

1.find the marginal revenue equation
2.find the marginal revenue for the production of 300,000,000 bushels
3. find the marginal revenue for the production of 450,000,000
Answer #1

1.find the marginal revenue equation (R'(x))

We have, revenue equation as R(x) = 0.737x2 + 1.4382x + 2.4811

Taking derivative wrt x, we get, R'(x) = d/dx(R(x))

R'(x) = d/dx(0.737x2 + 1.4382x + 2.4811)

R'(x) = 2*0.0737x + 1.4382

Thus, R'(x) = 0.1474x + 1.4382

2.find the marginal revenue for the production of 300,000,000 bushels

Given x is in hundreds of million dollars

thus, here x = 3

Puting x=3 in R'(x) we get

R'(3) = 0.1474*3 + 1.4382 = 1.8804

3. find the marginal revenue for the production of 450,000,000

Thus, here x = 4.5

Puting x=4.5 in R'(x) we get

R'(4.5) = 0.1474*4.5 + 1.4382 = 2.1015